OpenStudy (anonymous):

An easy one. Let P be a six degree polynomial. If P(1) = 1, P(2) = 2 ........till P(6) = 6. Find P(7) = ?

OpenStudy (anonymous):

And NO. Its not 7.

OpenStudy (anonymous):

The answer can be any number \[ P[(x)= \frac{x^6}{360}-\frac{7 x^5}{120}+\frac{35 x^4}{72}-\frac{49 x^3}{24}+\frac{203 x^2}{45}-\frac{39 x}{10}+2 \] P[7]=9 Here is another one \[\frac{x^6}{180}-\frac{7 x^5}{60}+\frac{35 x^4}{36}-\frac{49 x^3}{12}+\frac{406 x^2}{45}-\frac{44 x}{5}+4 \] P[7] =11 All depends on how the constant term is defined.

OpenStudy (anonymous):

In your explanation, your expression isnt getting the conditions satisfied. P(1) = 1 P(2) = 2 P(3) = 3 P(4) = 4 P(5) = 5 P(6) = 6

OpenStudy (anonymous):

Yes it is. Which one is not?

OpenStudy (anonymous):

Here is the general polynomial in terms of the conatant term a0 \[ P(x)= \frac{\text{a0} x^6}{720}-\frac{7 \text{a0} x^5}{240}+\frac{35 \text{a0} x^4}{144}-\frac{49 \text{a0} x^3}{48}+\frac{203 \text{a0} x^2}{90}+\frac{1}{20} (20-49 \text{a0}) x+\text{a0} \] \[ P[7} =7+ a0 \]

OpenStudy (anonymous):

P(7) =7+ a0

OpenStudy (anonymous):

A polynomial of degree 6 depends on 7 coefficients. You are given 6 conditions, so you can choose p(0) to be any thing and solve in temrs of a0. and P(7)= a0+7

OpenStudy (amistre64):

i think a matrix would help out

OpenStudy (amistre64):

we have 1 free variable making this infinite number of polys that fit as elias pointed out

OpenStudy (amistre64):

\[\begin{vmatrix}0\\0\\0\\0\\0\\1\\0\end{vmatrix}+k\begin{vmatrix}1/720\\-7/240\\35/144\\-49/48\\203/90\\-49/20\\1\end{vmatrix}\]

OpenStudy (anonymous):

@eliassaab That's really good. How did you think of those expressions btw?

OpenStudy (anonymous):

I mean my first answer was, \[P(x) = (x-1)(x-2)(x-3)(x-4)(x-5)(x-6) + x\]

OpenStudy (anonymous):

I solved 6 equations in six unknowns using a symbolic manipulator.

OpenStudy (anonymous):

That is one of the infinite possibilities. Probably the nicest one.

OpenStudy (anonymous):

Haha. True that. Good work. Thanks :)

OpenStudy (anonymous):

This would work too P(x)=K (x−1)(x−2)(x−3)(x−4)(x−5)(x−6)+x Where K is any constant.

OpenStudy (anonymous):

Yes. And that easily proves that it can have infinite permutations. Nicccce.